# Discovering Message Passing Hierarchies for Mesh-Based Physics Simulation

### v1

@@ D.2 Edge Enhancement
    Table 9: Results for different $K$ values in edge enhancement. $K=1$ means directly use the edges of selected nodes in the previous graph levels.
    RMSE-1 RMSE-50 RMSE-ALL
    $K=1$ 0.3296 54.22 100.1
    $K=2$ 0.3049 50.94 76.16
    $K=3$ 0.3380 52.71 86.84
    $K=4$ 0.3510 54.08 105.4
    Appendix E Stability Analysis
    Given the inherent randomness introduced by the Gumbel-Softmax sampling process in DiffSELECT, we evaluated the stability of DHMP by running the trained model on the test set in three independent trials. We report the mean and standard deviation of the prediction errors in Table 10. Despite the stochastic nature of the node selection process, the results show a very small standard deviation, demonstrating that DHMP reliably constructs stable and consistent dynamic hierarchies.
@@ Appendix E Stability Analysis
    This stability can be attributed to the DiffSELECT operation, where the node update module $\phi^{v}$ generates probabilities for retaining nodes in the next-level graph based on anisotropic aggregated edge features. The Gumbel-Softmax technique, coupled with temperature annealing, enables differentiable and stable node selection across hierarchy levels. As a result, the dynamic hierarchies are constructed in a manner that is not only consistent but also optimized for long-range dependencies.
    Moreover, the prediction errors from DHMP are significantly smaller than those of the baseline models, underscoring the robustness and reliability of the model, even with its dynamic node selection mechanism.
    Table 10: RMSE-1, -All (×$\times 10^{-2}$) for the trained DHMP across four Datasets with three independent tests.
    Model Cylinder Airfoil Flag Plate
    RMSE-1 0.1506 ±3.6E-4 36.27 ±5.7E-4 0.2741 ±2.4E-2 0.0263 ±5.6E-6
    RMSE-All 6.317 ±0.33 2018 ±130 68.66 ±2.9 1.327 ±0.002
    Appendix F Result Details
@@
    We report the mean and standard deviation for our model and baselines, calculated over three random seeds in training for Table 2. The results are shown in Table 11 and Table 12.
    Table 11: RMSE-1(×$\times 10^{-2}$) for our model and baselines from three different seeds.
    Model Cylinder Airfoil Flag Plate
    MGN 0.4046 ±1.08E-2 77.38 ±1.34E+1 0.4890 ±6.34E-2 0.0579 ±2.64E-3
    BSMS-GNN 0.2263 ±4.39E-2 71.69 ±1.41E+1 0.5080 ±0.48E-2 0.0632 ±14.3E-3
    Lino et al. 3.9352 ±11.3E-2 85.66 ±0.35E+1 0.9993 ±2.44E-2 0.0291 ±0.19E-3
    HCMT 0.9190 ±61.2E-2 48.62 ±0.51E+1 0.4013 ±1.76E-2 0.0295 ±3.45E-3
    DHMP 0.1568 ±0.94E-2 41.41 ±0.66E+1 0.3049 ±6.34E-2 0.0282 ±2.65E-3
    Table 12: RMSE-All(×$\times 10^{-2}$) for our model and baselines from three different seeds.
    Model Cylinder Airfoil Flag Plate
    MGN 59.78 ±2.00E+1 2816 ±1.99E+2 124.5 ±1.30E+1 3.982 ±1.14E-2
    BSMS-GNN 16.98 ±0.12E+1 2493 ±1.70E+2 168.1 ±0.65E+1 1.811 ±0.42E-2
    Lino et al. 27.60 ±0.86E+1 2080 ±0.39E+2 118.2 ±0.58E+1 2.090 ±13.2E-2
    HCMT 23.59±1.38E+1 3238 ±3.62E+2 90.32 ±0.50E+1 2.468±42.4E-2
    DHMP 6.571 ±0.06E+1 2002 ±1.02E+2 76.16 ±1.30E+1 1.296 ±1.14E-2
    Appendix G Computation Efficiency
@@
    We evaluated computational efficiency based on three criteria: training cost required to reach model convergence, inference time per step, and the total number of model parameters. A performance improvement threshold of less than 1% was used as the criterion for model convergence. The results are presented in Table 13.
@@
    Table 13: The detailed measurements of computation efficiency for DHMP and baseline models.
    Measurements Dataset MGN BSMS-GNN HCMT DHMP
    Training cost (hrs) Cylinder 35.26 37.11 80.60 35.96
    Airfoil 92.82 79.09 114.32 75.45
    Flag 28.12 18.27 66.70 17.14
    Plate 61.82 39.80 99.84 41.85
    Infer time/step (ms) Cylinder 17.35 16.55 79.52 21.79
    Airfoil 50.67 38.04 106.34 58.84
    Flag 16.15 17.18 85.87 26.33
    Plate 38.98 28.44 100.78 47.45
    #Parameter Cylinder 2.79M 2.05M 2.03M 2.66M
    Airfoil 2.79M 2.58M 2.03M 2.27M
    Flag 2.80M 2.06M 2.03M 2.67M
    Plate 2.80M 2.87M 2.03M 3.20M
    Appendix H Constructed Dynamic Hierarchies
    We visualize the constructed context-aware and temporally evolving hierarchies in Figure 8. We can see that the constructed hierarchies evolve as the input context changes and the evolving graph structures align with high-intensity regions.
    Figure 8: Row 1: the velocity field from the true data on CylinderFlow dataset. Row 2-6: the temporal difference of the velocity fields between adjacent time steps alongside the constructed coarser-level graphs.
    Appendix I Rollout Errors
@@
    Figure 910, and 11 showcase rollout error maps for the Airfoil, Flag, and DeformingPlate datasets. DHMP exhibits much lower rollout errors compared to the baseline.
@@
    Figure 9: Showcases of rollout prediction error maps on Airfoil dataset.
    Figure 10: Showcases of rollout prediction error maps on Flag dataset.
    Figure 11: Showcases of rollout prediction error maps on DeformingPlate dataset.

### v2

@@ C.3 Edge Enhancement
Table 10: Results for different values of $K$ in edge enhancement. Here, $K=1$ denotes directly using edges of selected nodes from previous graph levels. Training time and memory usage are measured with a batch size of $32$, while inference time and memory are evaluated with a batch size of $1$.
Training Infer
RMSE-1 (×$\times 10^{-2}$) RMSE-All (×$\times 10^{-2}$) Time (ms) vRAM (GBs) Time (ms) vRAM (GBs)
$K=1$ 0.3296 100.1 31.57 14.75 23.60 1.17
$K=2$ 0.3049 76.16 33.67 16.53 26.33 1.24
$K=3$ 0.3380 86.84 34.67 18.49 33.21 1.25
$K=4$ 0.3510 105.4 35.27 18.76 32.25 1.28
C.4 Hyperparameter Analyses on Number of Hierarchies
We conduct an ablation study to assess the impact of varying numbers of hierarchies on model performance. The results from the CylinderFlow dataset, illustrated in Figure 7, demonstrate that EvoMesh consistently outperforms BSMS-GNN across all tested numbers of graph hierarchies. Both models show improved performance with increased hierarchy depth up to $7$, indicating that deeper levels help capture more complex interactions and thus enhance accuracy. However, a slight performance decline is observed at level $9$, which may suggest the onset of overfitting. Overall, the dynamically learned hierarchies in EvoMesh are shown to be more effective compared to the predefined static hierarchies used in BSMS-GNN.
Figure 7: Model comparisons on different numbers of graph hierarchies on CylinderFlow dataset.
C.5 Effectiveness of Evolving Hierarchies
From Figure 4, by comparing EvoMesh vs. M2 and M4 vs. M1, we observe the advantages of learning adaptive and temporally evolving hierarchical graph structures. These results highlight the significance of adaptively modeling interactions in context-dependent graphs. To better understand how EvoMesh constructs dynamic hierarchies, we visualize the distribution of nodes with the top $10\%$ prediction errors in Figure 8(a). Accordingly in Figure 8(b), we observe that EvoMesh retains a higher proportion of “challenging” nodes in the coarser message passing levels, enabling our model to capture multi-scale dependencies more effectively, especially in areas where finer message passing levels struggle. In contrast, the predefined static hierarchies in the Bi-stride pooling baseline are data-independent and may inevitably overlook modeling long-range relations surrounding these pivotal nodes, even though they typically present higher errors than those in EvoMesh.
(a) Error distribution for BSMS-GNN and EvoMesh
(b) Ratios of challenging nodes
Figure 8: (a) Error maps, where nodes with the top $10\%$ of errors in each model’s predictions are marked in yellow and referred to as “challenging nodes”. (b) EvoMesh retains more challenging nodes in coarser graph hierarchies to capture multi-scale dependencies.
C.6 Stability Analysis
@@
Given the inherent randomness introduced by the Gumbel-Softmax sampling process in DiffSELECT, we evaluated the stability of EvoMesh by running the trained model on the test set in three independent trials. We report the mean and standard deviation of the prediction errors in Table 11. Despite the stochastic nature of the node selection process, the results show a very small standard deviation, demonstrating that EvoMesh reliably constructs stable and consistent dynamic hierarchies. This stability can be attributed to the DiffSELECT operation, where the node update module $\phi^{v}$ generates probabilities for retaining nodes in the next-level graph based on anisotropic aggregated edge features. The Gumbel-Softmax technique, coupled with temperature annealing, enables differentiable and stable node selection across hierarchy levels. As a result, the dynamic hierarchies are constructed in a manner that is not only consistent but also optimized for long-range dependencies. Moreover, the prediction errors from EvoMesh are significantly smaller than those of the baseline models, underscoring the robustness and reliability of the model, even with its dynamic node selection mechanism.
Table 11: Evaluation of EvoMesh with three independent tests.
Cylinder Airfoil Flag Plate
RMSE-1 (×$\times 10^{-2}$) 0.1506 ±3.6E-4 36.27 ±5.7E-4 0.2741 ±2.4E-2 0.0263 ±5.6E-6
RMSE-All (×$\times 10^{-2}$) 6.317 ±0.33 2018 ±130 68.66 ±2.9 1.327 ±0.002
C.7 Full Results over Multiple Training Seeds
@@
In Table 2 in the main manuscript, we report the mean results calculated over three random seeds. In Table 12, we provide full comparisons between our model and the baseline models, including standard deviations.
Table 12: Full quantitative results over three training seeds.
Cylinder Airfoil Flag Plate
RMSE-1 (×$\times 10^{-2}$)
MGN (2021) 0.3046 ±1.08E-2 77.38 ±1.34E+1 0.3459 ±6.34E-2 0.0579 ±2.64E-3
Lino et al. (2022) 3.9352 ±11.3E-2 85.66 ±0.35E+1 0.9993 ±2.44E-2 0.0291 ±0.19E-3
BSMS-GNN (2023) 0.2263 ±4.39E-2 71.69 ±1.41E+1 0.5080 ±0.48E-2 0.0632 ±14.3E-3
Eagle (2023) 0.1733 ±3.02E-2 0.385 ±1.03E+1 51.55 ±0.87E-2 0.0392 ±0.52E-3
HCMT (2024) 0.9190 ±61.2E-2 48.62 ±0.51E+1 0.4013 ±1.76E-2 0.0295 ±3.45E-3
EvoMesh 0.1568 ±0.94E-2 41.41 ±0.66E+1 0.3049 ±6.34E-2 0.0282 ±2.65E-3
RMSE-All (×$\times 10^{-2}$)
MGN (2021) 59.78 ±2.00E+1 2816 ±1.99E+2 115.3 ±1.30E+1 3.982 ±1.14E-2
Lino et al. (2022) 27.60 ±0.86E+1 2080 ±0.39E+2 118.2 ±0.58E+1 2.090 ±13.2E-2
BSMS-GNN (2023) 16.98 ±0.12E+1 2493 ±1.70E+2 168.1 ±0.65E+1 1.811 ±0.42E-2
Eagle (2023) 20.05 ±0.67E+1 2344 ±2.11E+2 127.7 ±0.88E+1 7.797 ±2.35E-2
HCMT (2024) 23.59 ±1.38E+1 3238 ±3.62E+2 90.32 ±0.50E+1 2.468 ±42.4E-2
EvoMesh 6.571 ±0.06E+1 2002 ±1.02E+2 76.16 ±1.30E+1 1.296 ±1.14E-2
C.8 Computation Efficiency
@@
We evaluate computational efficiency based on four criteria: training hours, inference time per step, and the total number of model parameters. The results are presented in Table 13.
@@
Table 13: The detailed measurements of computation efficiency for EvoMesh and baseline models.
      Measurements       Dataset       BSMS-GNN       HCMT       EvoMesh
      Training cost (hrs)       Cylinder       37.11       80.60       35.96
      Airfoil       79.09       114.32       75.45
      Flag       18.27       66.70       17.14
      Plate       39.80       99.84       41.85
      Infer time/step (ms)       Cylinder       16.55       79.52       21.79
      Airfoil       38.04       106.34       58.84
      Flag       17.18       85.87       26.33
      Plate       28.44       100.78       47.45
      # Parameter       Cylinder       2.05M       2.03M       2.66M
      Airfoil       2.58M       2.03M       2.27M
      Flag       2.06M       2.03M       2.67M
      Plate       2.87M       2.03M       3.20M
C.9 Rollout Showcases
@@
Figures 9 showcase rollout error maps for the Airfoil, Flag, and DeformingPlate datasets. EvoMesh exhibits much lower rollout errors than the baseline models.
(a) Airfoil
(b) Flag
(c) DeformingPlate
@@
Figure 9: Showcases of rollout prediction error maps on Airfoil, Flag and DeformingPlate dataset.
C.10 Constructed Dynamic Hierarchies
We visualize the constructed context-aware and temporally evolving hierarchies in Figure 10. We can see that the constructed hierarchies evolve as the input context changes and the evolving graph structures align with high-intensity regions. More visualizations of the evolution of the graph structure across the sequence are included in the supplementary material.
Figure 10: Row 1: The velocity field from the true data on the CylinderFlow dataset. Row 2-6: The temporal difference of the velocity fields between adjacent time steps alongside the constructed coarser-level graphs.
Appendix D Baseline Details
We compare EvoMesh with following competitive baselines: (1) MGN (Pfaff et al., 2021) which performs multiple message passing on the input high-resolution mesh topology; (2) Lino et al.(Lino et al., 2022), which uses manually set grid resolutions and spatial proximity for graph pooling; (3) BSMS-GNN (Cao et al., 2023), which uses predefined bi-stride pooling prior as preprocessing to generate static hierarchies on same mesh topology; (4) Eagle (Yu et al., 2024), which uses a two-scale hierarchical message-passing approach, downscaling mesh resolution via geometric clustering of mesh positions; (5) HCMT (Yu et al., 2024), which uses Delauny triangulation based on bi-stride nodes and adopt attention mechanism to enable non-uniform feature propagation. The architecture details of the compared models are as follows:
  • MGN.  In MGN, we use $15$ message passing steps in all datasets. The encoder, decoder, node update function, and edge update function are configured in the same way as in our model.
  • Lino et al.  We use the four-scale GNN structure proposed in the work of Lino et al. (2022). The edge length of the smallest cell for each dataset is $1/10$ of the average scene size, with each lower scale doubling in size. We follow its original paper to use $4$ message passing steps at the top and bottom levels and two for the others.
  • BSMS-GNN.  We use the same number of graph hierarchies in EvoMesh and as in BSMS-GNN. We use the minimum average distance as the seeding heuristic for the BFS search recommended in its original paper. The multi-level building is processed in one pass. The inter-level propagation uses the normalized node degree to convolve features from neighbors to central nodes. The encoder, decoder, node update function, and edge update function are set up the same way as in our model. We perform one message passing step at each graph level.
  • Eagle.  We use the same-size KMeans algorithm in the preprocessing step with a cluster size of 10. The encoder consists of 4 stacked GNN layers. The graph pooling module comprises a single-layer gated recurrent unit followed by a single-layer MLP. The hidden dimension is set to 128 across all datasets. The attention module includes 4 sequential attention blocks, each with a single attention head. A final layer normalization is applied after the last attention block. The decoder shares the same architecture as that used in EvoMesh.
  • HCMT.  The hidden dimension and the number of attention heads in the HCMT block are set to $128$ and $4$, respectively. We use the same number of hierarchies as in EvoMesh. For the Cylinder and Airfoil datasets, due to the presence of hollow sections in the mesh, we do not apply Delaunay triangulation for remeshing. Instead, we use edge connections generated through bi-stride pooling.Like in EvoMesh, we use a single message passing step at each graph level.
    Notably, the node encoder, decoder, node update function, and edge update function of MGN, Lino et al., BSMS-GNN, HCMT and Eagle have the same network architecture as those in EvoMesh. To reduce the number of network parameters, we avoid separately encoding the edge offset $\mathbf{e}_{ij}$. Instead, we concatenate it with the node latents and use this combined input for the edge update function to compute $\mathbf{\hat{e}_{ij}}$.
    All models are trained using the Adam optimizer for $1M$ steps, with an exponential learning rate decay from $10^{-4}$ to $10^{-6}$ and a decay rate of $\gamma=0.79$ from the first $500K$ steps. The batch size is set to $32$.